OpenStudy (anonymous):

h(x)=underroot3x-12 find the domain

OpenStudy (phi):

\[ h(x)= \sqrt{3x-12} \] So the previous problem was: NO DIVIDE BY ZERO. this one is : NO SQUARE ROOTS OF A NEGATIVE NUMBER

OpenStudy (phi):

the "bad x's" will make 3x-12 < 0

OpenStudy (anonymous):

greater thn 0? how

OpenStudy (anonymous):

less then i mean

OpenStudy (phi):

You need to solve relations: 3x-12 < 0 Have you seen problems like this (but never learned) or you never saw it before?

OpenStudy (anonymous):

never seen i guess

OpenStudy (phi):

could you find x if it were an equation 3x-12 = 0

OpenStudy (anonymous):

x=4

OpenStudy (phi):

if you change the = to < that is the answer x< 4 relations are almost the same as equations EXCEPT IF YOU DIVIDE BY A NEGATIVE NUMBER

OpenStudy (phi):

so to find the good x's we want 3x-12>0 can you solve this?

OpenStudy (anonymous):

x>4

OpenStudy (phi):

that is your domain (all good x's)

OpenStudy (anonymous):

so x>4 which means x not equal to 0

OpenStudy (phi):

wait. we want 3x-12≥ 0 x ≥=4 because x=4 is allowed.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

getting it now

OpenStudy (anonymous):

what abut the under root?

OpenStudy (phi):

x≥4 means x greater than or equal to 4 numbers like 4, 4.1, 5, 5.3444 , 37, etc, etc...

OpenStudy (anonymous):

so my answer is x/x>4 .

OpenStudy (phi):

You have found all the x's that give a 0 or positive number under the root, so f(x) will have a valid number. All the numbers that f(x) can be is the RANGE of f(x)

OpenStudy (anonymous):

I need to find the x and then write it in set builder notation

OpenStudy (phi):

almost: X where X≥4 (X>4 would exclude 4, but 4 is ok... try it)

OpenStudy (anonymous):

and the number is 4 we found out by calculating the equation.

OpenStudy (phi):

\[ \{ x | x\in R, x≥4 \} \]

OpenStudy (phi):

and the number is 4 we found out by calculating the equation. we found 4 by insisting that 4x-12 ≥ 0 (the stuff inside the root must not be negative)

OpenStudy (anonymous):

and what we did with the sq root?

OpenStudy (anonymous):

ok Now I am getting the concept

OpenStudy (phi):

The question was find the domain that means what are all the good x's ? square root is used to find f(x), not the good x's

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here is the next situation in next question

OpenStudy (anonymous):

if it is 4 / under root x-9?

OpenStudy (phi):

looks like we have to use both rules: NO DIVIDE BY ZERO, NO SQUARE ROOTS OF NEGATIVE NUMBERS \[ f(x)=\frac{4}{\sqrt{x-9}} \] can you write the relation for all "good x's"?

OpenStudy (anonymous):

x=4/9

OpenStudy (anonymous):

x/x>4/9

OpenStudy (phi):

you want to look just at the stuff under the root. It must be positive. x>4/9 would mean 1 should be good. (1 is bigger than 4/9, right?) by 1-9 is -8, and no negatives allowed inside the square root. so just look at x-9

OpenStudy (anonymous):

x/x>9 onlyy and for get about 4?

OpenStudy (phi):

remember what you are looking for: negatives inside the root. the 4 up top won't affect anything. also, make sure you do not get 0: Let's check if x were 9 the bottom would be sqrt(0) =0 BAD. so x>9 is the answer.

OpenStudy (anonymous):

and if under root t-4/3t-21. so in this case we will look only on the top under root one>

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and for get the bottom one?

OpenStudy (phi):

better post that. It is complicated.

OpenStudy (anonymous):

ok